OpenStudy (anonymous):

The width of a rectangle is 7 less than twice its length. If the area of the rectangle is 113 cm^2, what is the length of the diagonal? i did x(2x-7)=113 did the quadformula and it didnt work :\

OpenStudy (anonymous):

the quad formula will give you the length of the rectangle. from this you can get the width. Once you have the length and the width, you get the length of the diagonal by using the pythagorean theorem.

jhonyy9 (jhonyy9):

2x^2 -7x -113=0 x_1,_2=(7+/- sqrt(49+904))/4 =(7+/- sqrt953)/4 =

OpenStudy (anonymous):

Yeah I did that but you end up getting decimals and it doesnt work in the hw site for some reason.... i got 9.47 for the length and 11.94 for the width based on the equation. Idk if I did something wrong

OpenStudy (anonymous):

hmm ill check your results

jhonyy9 (jhonyy9):

so if sqrt953 =30,8 -so continue from this

OpenStudy (anonymous):

length is 9.47, width is 11.92. Length of the diagonal is 15.2

OpenStudy (anonymous):

nope 15.2 doesnt work either :\

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

stupid question but...it wants the length of the diagonal is cm, right?

OpenStudy (anonymous):

in*

OpenStudy (anonymous):

Yeah im assuming

OpenStudy (anonymous):

it lists _____cm though

jhonyy9 (jhonyy9):

x=37,8/4 =9,4 w=2l-7=18,8-7=11,8 diagonal d=sqrt(139.24+88,36)=sqrt227,6 =15,08

OpenStudy (anonymous):

not sure what the issue with your hw site is; make sure you carry all your digits forward prior to rounding and other than that, your correct

OpenStudy (anonymous):

carrying all the digits worked thank you.

OpenStudy (anonymous):

ok good.. :)